OpenStudy (anonymous):

prove: sinθ - sinθ * cos^2 θ = sin^3 θ, show all work please!

OpenStudy (anonymous):

@hartnn Im sorry, theres one more for my guide i missed XC

hartnn (hartnn):

did u know ? \(\Large \sin^2 \theta +\cos^2\theta =1 \) ?

OpenStudy (anonymous):

nope, im 100% unsure of how this question works

hartnn (hartnn):

so now you know :) isolate cos^2 theta from that formula

OpenStudy (anonymous):

sin^2θ=1?

OpenStudy (anonymous):

wait wait

OpenStudy (anonymous):

Nope, I'm lost ._.

hartnn (hartnn):

\(1-\cos^2 \theta =1- (1-\sin^2 \theta ) = \sin^2 \theta \) see if you get this ?

OpenStudy (anonymous):

Ok i understand that

hartnn (hartnn):

sinθ - sinθ * cos^2 θ factor out sinθ what do u get?

OpenStudy (anonymous):

sin^3(θ) ?

hartnn (hartnn):

left side = sinθ - sinθ * cos^2 θ = sinθ ( 1- cos^2 θ) = sin(θ) * sin^2(θ) =sin^3(θ) = right side hence proved

OpenStudy (anonymous):

You make this so easy ._. Before you math was just mumbo jumbo but after you explain it it seems so obvious!

hartnn (hartnn):

thats a nice compliment! you're welcome ^_^